OpenStudy (wannabegurl):

How can I tell weather I add or subtract to find a the missing length of a triangle , using the polygorean theorem

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):



OpenStudy (misty1212):

\[a^2+b^2=c^2\\ c^2-b^2=a^2\] for example

OpenStudy (irishboy123):

that's *cool squared* !!! "polygorean theorem" like off Star Trek?!?! corresponds with our Terrestrian version. AFAIK.